OpenStudy (anonymous):

the legnth of a rectangle is 3cm greater then its width.the primeter is 24 cm.find the dimensions of the rectangle. A. 3cm,9cm B. 4.5cm,7.5cm C. 7.5cm, 10.5cm D. 10.5cm,13.5cm

OpenStudy (anonymous):

@triciaal one more plz

OpenStudy (anonymous):

ion even kno doe plz help

jhonyy9 (jhonyy9):

a is right sure 3+3+9+9=24

OpenStudy (anonymous):

i just got dat 1 wrong doe

jhonyy9 (jhonyy9):

sorry then 3 time greater will be right A. in this case that with 3 cm is greater so then B is right

OpenStudy (anonymous):

I don't likr msth

jhonyy9 (jhonyy9):

4,5+4,5=9 7,5+7,5=15 9+15=24

jhonyy9 (jhonyy9):

7,5 is with 3 cm greater then 4,5 yes ?

jhonyy9 (jhonyy9):

and in case of A. 9 is 3 time greater then 3 right ?

jhonyy9 (jhonyy9):

so i think this is more understandably now yes ?

jhonyy9 (jhonyy9):

hope so was my pleasure bye

OpenStudy (triciaal):